OpenStudy (anonymous):

As the principal quantum number increases, the energy difference between successive energy levels decreases. Why?

OpenStudy (abb0t):

As n increases for a given atom, so does the average distance of an electron from the nucleus

OpenStudy (anonymous):

but how does that explain why the energy difference between n=1 and n=2 would be greater than n=4 and n=5?

OpenStudy (anitaasks):

The energy of an orbital is given by the formula, energy of orbital= -2.18E-18(Z^2/n^2) so as n increases, the energy is smaller

OpenStudy (aaronq):

The principle quantum number relates to the distance of an electron to the nucleus. if you model an atom with the nucleus (protons) as a point charge, you can see that electric field of the atom decays with as an inverse square of the radius (\(E=\dfrac{1}{r^2}\)). To illustrate how the energy difference between successive energy levels decreases, compare the differences of two distances from the nucleus, 0.1 and 0.2 vs 0.4 and 0.5 \(\dfrac{1}{(0.1)^2}=100\) \(\dfrac{1}{(0.2)^2}=25\) difference of 75 \(\dfrac{1}{(0.4)^2}=6.25\) \(\dfrac{1}{(0.5)^2}=4\) difference of 2.25